OpenStudy (anonymous):

In how many ways can the letters of the word STORAGE be arranged using only 4 letters at a time?

OpenStudy (anonymous):

like rearranging it into a new word like rage?

OpenStudy (anonymous):

Yes order does not matter

OpenStudy (anonymous):

it depends on what language lol jk and i think about five or six

OpenStudy (anonymous):

5040 28 210 840

OpenStudy (welshfella):

7C4 - hte number of combinations of 4 in 7

OpenStudy (anonymous):

Ooo never mind my answer..... And 28 because 4 times 7 equals 28

OpenStudy (anonymous):

So it would be B?

OpenStudy (anonymous):

yes B would be correct

OpenStudy (anonymous):

Thank you!

OpenStudy (welshfella):

if the order of the letters in the 4 do matter its 210

OpenStudy (anonymous):

Oh okay. Thanks! The question doesn't really tell me if it does or doesn't matter? What would be the most likely?

OpenStudy (anonymous):

no not really it would still always be 7*4=28 so 28 is the most likely answer

OpenStudy (anonymous):

Okay Thank you all for your help!

OpenStudy (welshfella):

i'm not sure its 28 because the number of combinations of 4 from 7 = 7*6*5/ 6 = 35 maybe theres a typo somewhere

OpenStudy (welshfella):

yw